A loan of ​$4908 borrowed today is to be repaid in three equal installments due in two years​, three ​years, and five dash and dash a dash half ​years, respectively. What is the size of the equal installments if money is worth 9.6 % compounded monthly question mark
Answer #1

Loan Amount = $4,908

Interest rate p.a = 9.6% compounded monthly

Interest rate p.m = 9.6%/12 = 0.8%

Months in 2 years = 2*12 =24

Months in 3 years= 3*12 = 36

Months in 5.5 years = 5.5*12 = 66

Let's assume equal size of installment is $X.

We can calculate value of X with following equation -

Thus, The equal installment would be $2,264.27

Hope this will help, please do comment if you need any further explanation. Your feedback would be appreciated.